# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 1.2 Outline of the proof
Coming back to the examples in rank two, for notational convenience let $\langle X_{v}\rangle$ denote the corresponding set of $B$-modules, $v$ running over the Weyl group $W$. Let ${\mathrm{rep}}(B)$ be the category of finitely generated $B$-modules. The first key observation is that in each type in rank two the $B$-modules $X_{v}$ have a relative Schubert filtration by modules $Q(\lambda)$ and that the dual $B$-modules $X_{v}^{*}$ have an excellent filtration by modules $P(\lambda)$; the second key observation is that the $X_{v}$’s involve in several ways the Steinberg weights $e_{v}$ known from the Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ of the representation ring $R(B)={\sf K}_{0}({\mathrm{rep}}(B))$ as a module over the representation ring $R(G)$, see Section 3.2.
Now it is the Steinberg basis and combinatorics related to it that take the central stage. In the set of modules $Q(\lambda)$ that occur in the filtration of $X_{v}$, the weight $\lambda$ of the $B$-socle of $Q(\lambda)$ is always a Steinberg weight $e_{w}$ with $w\geq v$ in the Bruhat order on $W$. Moreover, the modules $P(-\lambda)^{*}$ that occur in the other filtration of $X_{v}$ are of the form $\nabla_{\mu}\otimes P(-e_{w})^{*}$ with $\mu$ dominant, $w\leq v$ in the Bruhat order. We show in Theorem 8.16 that one still gets a basis of $R(B)$ when replacing a few ${\sf k}_{e_{v}}$’s with $Q(e_{v})$ or $P(-e_{v})^{*}$, see Section 8.
The characters of the modules $Q(e_{v})$ and $P(-e_{w})$ may be computed with Demazure operators, and with computer assisted computations in rank three we have found a close relation in the $T$-equivariant ${\sf K}$-group ${\sf K}_{T}(G/B)$ between the classes of the $\mathcal{L}(P(-e_{w}))$ and the Schubert basis of the free $R(T)$-module ${\sf K}_{T}(G/B)$. We show in Theorem 3.1 that the transition matrix between the basis $\{\mathcal{L}(P(-e_{w}))\}_{w\in W}$ and the Schubert basis of ${\sf K}_{T}(G/B)$ is always invertible and triangular, up to a permutation of rows and columns. Similarly, we show there is a triangular transition matrix comparing the basis $\{\mathcal{L}(Q(e_{v}))\}_{v\in W}$ with the opposite Bruhat cell basis, see Section 5.
This triangularity turns out to be the key to the desired ${\mathrm{Ext}}$-vanishing required for exceptional collections. Precisely, much of the required coherent cohomology vanishing is assured by the results from [47], e.g. Theorems 6.2 and 6.4. They allow to reduce the problem to Euler characteristic computations. Then the crucial vanishing results are provided by the results of Section 5. The precise statement is in Theorem 6.5.
Our way to construct full exceptional collections on flag varieties $G/B$ goes through the study of the derived category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a category which is linear over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ (a $G$-linear category for short), see Section 8.2. Note that the action of the latter category on the former is given by the restriction functor ${\mathrm{res}}_{B}^{G}$, cf. the aforementioned quote from Donkin. We first construct semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $G$-linear category into pieces, each equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ (a $G$-linear semiorthogonal decomposition for short, see Definition 20). Sections 8 and 9 are devoted to constructing objects $X_{p}$, $p\in W$ labelled by the elements of the Weyl group; these objects $X_{p}$, $p\in W$ will serve as the basis for constructing $G$-semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. This requires choosing a total order $\prec$ on $W$ that refines the Bruhat order. This phase itself breaks up into two steps: given an element $p\in W$, the first step constructs a $G$-semiorthogonal decomposition (called “cut at $p$”, see Section 9.1) of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ into two subcategories
@@
$\langle\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}),\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})\;\rangle.$ (2)
In turn, to establish (2) one needs the crucial vanishing result of Corollary 6.7 that ensures semiorthogonality of those subcategories, and one also needs Theorem 8.15 which says that the triangulated hull of the union of the two categories in (2) is the whole of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. The latter property is a corollary of the generation property saying that the $B$-modules in (2) generate the abelian category ${\mathrm{rep}}(B)$. This relies on Weyl group’s combinatorics and is explained in Section 8.6. As a byproduct of the results of Section 8, we obtain an alternative proof of a theorem of Steinberg [44].
@@
Using $G$-linear semiorthogonal decompositions as in (2), we can finally construct the objects $X_{p}$, $p\in W$ of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. In a nutshell, the procedure is as follows. Given an element $p\in W$, associated to it are the $B$-modules $P(-e_{p})^{\ast}$ and $Q(e_{p})$ where $e_{p}$ is the Steinberg weight associated with $p$. We define $X_{p}$ to be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. Similarly, we define $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ into $\mathcal{D}$. Section 10 proves that $X_{p}$ and $Y_{p}$ are isomorphic for any $p\in W$, and that the objects $X_{p}$, $p\in W$ are $B$-exceptional, i.e. $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))}(X_{p},X_{p})={\sf k}$.
@@
Now each object $X_{p}$, $p\in W$ gives rise to a $G$-linear functor $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Cohomological vanishing statements from Section 6 combined with Theorem 11.1 prove that each $G$-linear functor $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ is full and faithful. Results of Section 8.2 then allow to establish that the collection of full triangulated subcategories ${\rm Im}(\Phi_{p})\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$, $p\in W$, forms a $G$-linear semiorthogonal decomposition of the category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ with respect to the chosen total order $\prec$ on $W$. This is the statement of Theorem 11.5.
@@
The full exceptional collection in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ is now obtained as a consequence of Theorem 11.5; this is taken care of in Theorem 12.1 of Section 12. Specifically, we define objects ${\mathcal{X}}_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$ as the images of $X_{p}$, $p\in W$ under the composition of the sheafification functor $\mathcal{L}$ with the forgetful functor ${\sf For}:{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}}(\mathop{\sf{Coh}}\nolimits^{G}(G/B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(G/B))$. The required cohomological vanishing for the objects ${\mathcal{X}}_{p}$, $p\in W$ is in fact provided by what has been previously done in Theorem 11.5, while the generating property is assured by the results of Section 8. For the details, see Theorem 12.1.
The appearance of $G$-equivariant vector bundles in the full exceptional collection is predicted by [39, Lemma 2.2]. We also knew from the rank two examples that we had to look for ${\mathcal{X}}_{v}$ inside the intersection of the hull of $\{\mathcal{L}(P(-e_{v})^{*})\}_{v\preceq p}$ with the hull of $\{\mathcal{L}(Q(e_{v}))\}_{v\succeq p}$.
In section 13 we extend our results to the case of a parabolic subgroup $P\subset G$ with $B\subset P$. We do not need to refer explicitly to highest weight category structures on ${\mathrm{rep}}(P)$. Such structures on ${\mathrm{rep}}(P)$ are available by [21] and appear implicitly in [40] the results of which also bear on the quivers from Section 14.

### v3

@@ 1.2 Outline of the proof
In the next two paragraphs, we assume again that the group $\mathbf{G}$ is of rank two. We choose a total order $\prec$ on the corresponding Weyl group that refines the Bruhat order. For notational convenience, let $\langle X_{v}\rangle,v\in W$ denote the set of $\mathbf{B}$-modules that underlie the full exceptional collection of $\mathbf{G}$-equivariant vector bundles on $\mathbf{G}/\mathbf{B}$ ordered with respect to $\prec$. Recall that ${\mathrm{rep}}(\mathbf{B})$ denotes the category of finitely generated $\mathbf{B}$-modules over ${\sf k}$. A detailed inspection of the modules $\langle X_{v}\rangle,v\in W$ led to the two key observations that made it possible to see the pattern. The first one is that in each type in rank two the $\mathbf{B}$-modules $X_{v}$’s have relative Schubert filtrations by modules $Q(\lambda)$, while the ${\sf k}$-linear dual $\mathbf{B}$-modules $X_{v}^{*}$’s have excellent filtrations by modules $P(\lambda)$. The second observation is the appearance of the Steinberg weights $e_{v}\in X(\mathbf{T})$ in weight filtrations on the modules $X_{v}$’s; by [46], those weights form a basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ (the Steinberg basis) of the representation ring $R(\mathbf{B})={\sf K}_{0}({\mathrm{rep}}(\mathbf{B}))$ as a module over the representation ring $R(\mathbf{G})$, see Section 4.2.
More precisely, in the set of modules $Q(\lambda)$ that occur in the filtration of $X_{v}$’s, the weight $\lambda$ of the $\mathbf{B}$-socle of $Q(\lambda)$ is always a Steinberg weight $e_{w}$ with $w\geq v$ in the Bruhat order on $W$. On the other hand, the modules $P(-\lambda)^{*}$ that occur in the other filtration of $X_{v}$’s are of the form $\nabla_{\mu}\otimes P(-e_{w})^{*}$ where $\nabla_{\mu}$ is a dual Weyl module for a dominant weight $\mu\in X(\mathbf{T})$ and $w\leq v$ in the Bruhat order. As the corresponding coherent sheaves $\mathcal{L}(X_{v}),v\in W$ on $\mathbf{G}/\mathbf{B}$ form a full exceptional collection in rank two by Section 14, this imposes a semiorthogonality condition on the sheaves $\mathcal{L}(P(-e_{w})^{*}),w\in W$ and $\mathcal{L}(Q(e_{v})),v\in W$ with respect to the Bruhat order. Furthermore, the $\mathbf{B}$-socle of the module $Q(e_{v})$ is ${{\sf k}}_{e_{v}}$, while the same character ${{\sf k}}_{e_{v}}$ is also the $\mathbf{B}$-head of the module $P(-e_{v})^{*}$. These facts in rank two will serve as the basis for producing the sought-for semiorthogonal decompositions for flag varieties in arbitrary rank; it starts with Theorems 6.5 and 6.6 (the reader is also invited at this point to keep consulting Section 1.4 for an easier navigation in the key statements across the text). It is also at this stage that the interplay among three orders - one on the Weyl group (the Bruhat order) and the other two on $X(\mathbf{T})$ (the excellent and antipodal excellent orders) becomes crucial in our considerations.
It should be noted that an earlier paper [4] unveiled the role played by the Steinberg basis in the setting of exceptional collections. At the same time, it showed its limitations already for the group ${\bf G}_{2}$ if the Steinberg basis is taken literally. As our paper explains, the crucial missing ingredient was that the Steinberg weights should have been seen through the lens of the category ${\mathrm{rep}}(\mathbf{B})$ – that is as socles (resp., heads) of costandard (resp. standard) objects for the highest weight category structure on ${\mathrm{rep}}(\mathbf{B})$ given by the antipodal excellent order. The remaining parts of the costandard modules $Q(e_{v}),v\in W$ (resp., of the standard modules $P(-e_{v})^{*},v\in W$) that lie above the socle (resp., below the head) of their common Steinberg weight ${{\sf k}}_{e_{v}}$ are responsible for the eventual semiorthogonality properties.
Now, as we have the Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$, and the $\mathbf{B}$-modules $P(-e_{v})^{*}$ and $Q(e_{v})$ with $v\in W$ at our disposal for a general semisimple group $\mathbf{G}$, let us take a closer look at the semiorthogonality condition on $\mathcal{L}(P(-e_{w})^{*})$ and $\mathcal{L}(Q(e_{v}))$ and explain how we arrived at those. Precisely, this is about Theorems 6.5 and 6.6 (triangularity of the transition matrices at the $\sf K$-theoretical level). The preparatory technical work for those theorems is done in Section 5 and proceeds as follows. Let $\mathbf{B}^{+}$ be the opposite Borel subgroup to $\mathbf{B}$. The $\mathbf{T}$-equivariant $\sf K$-theory of a flag variety $\mathbf{G}/\mathbf{B}^{+}$ for a Borel subgroup $\mathbf{B}^{+}$ has two natural bases: one consisting of the classes $\{[\mathcal{O}_{X_{w}}],w\in W\}$ of the closures of $\mathbf{B}^{+}$-orbits on $\mathbf{G}/\mathbf{B}^{+}$ (the “Schubert basis”), and another consisting of the classes $\{[\mathcal{O}_{{X}^{w}}(-\partial{X}^{w})]\}_{w\in W}$ of the $\mathbf{B}^{-}$-orbits on $\mathbf{G}/\mathbf{B}^{+}$ for the opposite Borel subgroup $\mathbf{B}^{-}$ to $\mathbf{B}^{+}$ (the “opposite Bruhat cell basis”). The characters of the modules $Q(e_{v})$ and $P(-e_{w})$ can be computed with Demazure operators, and with computer assisted computations in rank three we have found a close relation in the $\mathbf{T}$-equivariant ${\sf K}$-group ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ between the classes $\{[\mathcal{L}(P(-e_{w}))],w\in W\}$ and the Schubert basis $\{[\mathcal{O}_{X_{w}}],w\in W\}$. More precisely, the $R(\mathbf{T})$-module ${\sf K}^{0}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ is equipped with a natural $R(\mathbf{T})$-valued paring, and a result from [28], which is attributed to Knutson in loc.cit., asserts that the Schubert and the opposite Bruhat cell bases are orthogonal to each other with respect to that pairing. Using this very important input, in Section 5 we show in Theorem 4.1 that the transition matrix between the basis $\{\mathcal{L}(P(-e_{w}))\}_{w\in W}$ and the Schubert basis of ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ is always invertible and triangular, up to a permutation of rows and columns. Similarly, the transition matrix from the basis $\{\mathcal{L}(Q(e_{v}))\}_{v\in W}$ to the opposite Bruhat cell basis is also invertible and triangular, up to a permutation of rows and columns. Finally, much of the cohomology vanishing required by the statement of Theorem 6.5 is assured by the results explained in [49], e.g. Theorems 6.2 and 6.4, which are based on [39]. These allow to reduce the proof of Theorem 6.5 to the Euler characteristic computations, and this is precisely what Section 5 does with the help of all the input explained above in this paragraph. The triangularity results of Theorems 6.5 and 6.6 turn out to be the key to the desired ${\mathrm{Ext}}$-vanishing required for exceptional collections. Arguably, Theorems 6.5 and 6.6 are the most important cohomological statements of our work.
Section 5, discussed in the previous paragraph, concerns with the structure of the $R(\mathbf{T})$-module ${\sf K}^{0}(\mathbf{G}/\mathbf{B})$. Similarly, the representation ring $R(\mathbf{B})={\sf K}_{0}({\mathrm{rep}}(\mathbf{B}))$ is a module over the representation ring $R(\mathbf{G})$; in both cases, the $R(\mathbf{T})$-module structure for the former module (resp., the $R(\mathbf{G})$-module structure for the latter) are essential for our purposes. More generally, the category ${\mathrm{rep}}(\mathbf{B})$ has a module structure over the base abelian category ${\mathrm{rep}}(\mathbf{G})$; that module structure is given by the restriction functor ${\mathrm{res}}_{\mathbf{B}}^{\mathbf{G}}$, cf. the aforementioned quote from Donkin. Passing to derived categories, the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ receives a module structure over the base category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ (since that module structure comes from a monoidal action at the level of abelian categories, it doesn’t require further compatibilities). In this framework, thanks to Theorems 6.5 and 6.6 the basic examples of Section 14 can be reformulated as saying that for a group $\mathbf{G}$ of rank two and a Borel subgroup $\mathbf{B}\subset\mathbf{G}$, there are semi-orthogonal decompositions of the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ that are linear over $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ ($\mathbf{G}$-linear semiorthogonal decompositions for short, see Definition 8.8) with each semiorthogonal component being equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
Given a base scheme $S$, there are robust notions of $S$-linear triangulated categories and of $S$-linear semiorthogonal decompositions of a given $S$-linear triangulated category, [36]. The notions of a $\mathbf{G}$-linear triangulated category and of a $\mathbf{G}$-linear semiorthogonal decomposition are in complete parallel and could be put on an equal footing with those from loc.cit. if we used the language of (quotient) stacks that we deliberately avoid. For that reason, we recall in Sections 8.2 and 8.3 all the necessary definitions and statements concerning $\mathbf{G}$-linear triangulated categories in the classical language of equivariant (quasi)-coherent sheaves.
Having set up the necessary framework, we proceed to constructing $\mathbf{G}$-linear semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ in arbitrary rank. For that we need to choose a total order $\prec$ on the Weyl group $W$ that refines the Bruhat order. For any element $p\in W$ of the Weyl group we will first construct in Section 9.1 a $\mathbf{G}$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ (called “cut at $p$”) into two admissible subcategories
@@
$\langle\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}}),\;{\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}})\;\rangle,$ (3)
Here ${\mathrm{hull}}(-)$ denotes the triangulated envelope of a given set of objects, and $\nabla_{\lambda}$ are dual Weyl modules. Strikingly, the $\mathbf{B}$-modules $P(-e_{v})^{*},v\in W$ and $Q(e_{v}),v\in W$ will serve for both semiorthogonality condition in (3), which will follow from Theorem 6.6 and Corollary 6.7, and for the generating property. Precisely, in Theorem 8.19 we prove that the union of the two triangulated hulls in (3) is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. That proof, which takes Section 8.6, can be considered as a categorical upgrade of [3, Theorem 2] which is essentially the same statement as in Theorem 8.19, but at the $\sf K$-theory level. The reader may prefer to consult [3, Theorem 2 and Section 4] as a starting point before proceeding to Section 8.6. As a byproduct of the results of Section 8, we obtain an alternative proof and categorification of a theorem of Steinberg [46], see Theorem 13.11, which is similar to [3, Theorem 2] of Ananyevskiy.
@@
Semiorthogonal decompositions (3) are the core part of our paper as they allow to construct objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ that will eventually give rise to full exceptional collections on $\mathbf{G}/\mathbf{B}$. More precisely, for an element $p\in W$ we define in Section 9.2 two objects $X_{p}$ and $Y_{p}$ of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Specifically, given the semiorthogonal decomposition (3) associated to $p\in W$, we define an object $X_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ to be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(\mathbf{T})_{+}}$ into $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Similarly, we define $Y_{p}$ be the image of $Q(e_{p})$ under the right adjoint of the inclusion of the triangulated hull of $\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}}$ into $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. The reason of defining the objects $X_{p}$ and $Y_{p}$ as above was prompted by the insight coming from the examples in rank two: the exceptional objects ${\mathcal{X}}_{v}$ on $\mathbf{G}/\mathbf{B}$ for rank two groups lie inside the intersection of the hull of $\{\mathcal{L}(P(-e_{v})^{*})\}_{v\preceq p}$ with the hull of $\{\mathcal{L}(Q(e_{v}))\}_{v\succeq p}$.
The final step of this core part is showing that for any $p\in W$ the objects $X_{p}$ and $Y_{p}$ are isomorphic. This is done in Sections 9.3 and 10. To this end, we need a refinement of our categorification of Steinberg’s theorem, which is Theorem 10.1. This refinement, implicit in [3], shows that the Steinberg weights serve as a curious meeting ground of the Bruhat order and the antipodal excellent order $<_{a}$ on the weight lattice $X(\mathbf{T})$ from Definition 2.5 (compare also Remark 10.3). At this stage we learn from Lemma 9.5 that the objects $X_{p}$’s we have constructed are $\mathbf{B}$-exceptional, i.e. $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))}(X_{p},X_{p})={\sf k}$; that is, they are exceptional in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ considered as a $\sf k$-linear category.
@@
But now, thanks to Theorem 6.6 and to the very construction of objects $X_{p},p\in W$, there is much more to cohomological properties of the $X_{p}$’s than asserts Lemma 9.5: these objects turn out to be exceptional in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ considered as a $\mathbf{G}$-linear category, see Theorem 11.1 for the precise statement. Now, by Proposition 8.11 each object $X_{p}$, $p\in W$ gives rise to a $\mathbf{G}$-linear functor $\Phi_{p}\colon\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Cohomological vanishing statements from Section 6 combined with Theorem 11.1 prove that each $\mathbf{G}$-linear functor $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ is full and faithful. Results of Sections 8.2 and 8.3 then allow to establish that the collection of full triangulated subcategories ${\rm Im}(\Phi_{p})\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$, $p\in W$, forms a $\mathbf{G}$-linear semiorthogonal decomposition of the category $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ with respect to the chosen total order $\prec$ on $W$. This is the statement of Theorem 11.6.
@@
Now the proof of Theorem 12.1 goes as follows. We define objects ${\mathcal{X}}_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$ as the images of $X_{p}$, $p\in W$ from Section 9 under the composition of the sheafification functor $\mathcal{L}$ with the forgetful functor ${\sf For}:{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}}(\mathop{\sf{Coh}}\nolimits^{\mathbf{G}}(\mathbf{G}/\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$. Theorem 11.6 implies that the collection of objects ${\mathcal{X}}_{p},p\in W$ is an exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$, while the generating property of ${\mathcal{X}}_{p},p\in W$ is assured by the results of Section 8. Remark 12.2 draws a parallel of Theorem 12.1 with a base change type of statement for semiorthogonal decompositions, [36].
Section 13 treats the case of a parabolic subgroup $\mathbf{P}\supset\mathbf{B}$. Its main statements, which are Theorems 13.3 and 13.5, follow essentially the same path that has been set out in Theorems 11.6 and 12.1. For the reasons that are explained below in this paragraph, it is natural to expect that the objects ${\hat{\sf X}}_{v},{v\in W^{\mathbf{P}}}$ that give rise to semiorthogonal decompositions of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ as a $\mathbf{G}$-linear category (resp., the objects $\hat{\mathcal{X}}_{v}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P}),v\in W^{\mathbf{P}}$ giving full exceptional collections in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P})$) are contained among the objects ${\sf X}_{v},{v\in W}$ of Theorem 11.6 (resp., among the objects ${\mathcal{X}}_{v},v\in W$ of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B})$ of Theorem 13.5). One has therefore to recognize those objects among ${\sf X}_{v},{v\in W}$ (resp., among ${\mathcal{X}}_{v},v\in W$) that are obtained by the restriction functor ${\mathrm{res}}_{\mathbf{G}}^{\mathbf{P}}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ (resp., by the pullback $\pi_{\mathbf{P}}^{\ast}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{P})\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B})$ along the projection $\pi_{\mathbf{P}}:\mathbf{G}/\mathbf{B}\rightarrow\mathbf{G}/\mathbf{P}$). The fundamental fact that both functors ${\mathrm{res}}_{\mathbf{G}}^{\mathbf{P}}$ and $\pi_{\mathbf{P}}^{\ast}$ are $t$-exact and fully faithful on the respective derived categories makes it possible to recognize the sought-for exceptional objects on $\mathbf{G}/\mathbf{P}$ by applying the induction functor ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{P}}$ to appropriate objects ${\sf X}_{v},{v\in W}$ (resp., the pushforward $R{\pi_{\mathbf{P}}}_{\ast}$ to ${\mathcal{X}}_{v},v\in W$). It turns out the Steinberg weights $e_{v},v\in W^{\mathbf{P}}$ for a given parabolic $\mathbf{P}$ behave nicely with respect to the induction functor ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{P}}$ suggesting a natural parabolic analogue of the key $\mathbf{B}$-modules from Section 2. The cohomological properties of those parabolic analogues are given by Theorem 13.16, a parabolic counterpart of Theorem 6.6. That allows to further apply the arguments of Sections 8 and 9 in the parabolic case obtaining Theorems 13.3 and 13.5.